I have a 67% average right now. My exam tomorrow counts for 40% of my final grade. what do i have to get on the exam in order to get my final grade to 70%
Answer by TimothyLamb(4379)

---
x = final exam grade
0.6*67 + 0.4x = 70
0.4x = 70 - 0.6*67
x = (70 - 0.6*67)/0.4
x = 74.5
---
answer:
must get 74.5% on the final exam
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php